OpenStudy (anonymous):

you bought a new tv price at $500 including tax. you paid nothing down and will make 24 payments of $25 each. how much interest were you charged

OpenStudy (anonymous):

You could yust Mathway.com

OpenStudy (anonymous):

i don't cheat my way to success. i want it well earned. i'm currently doing the problem.

OpenStudy (anonymous):

any assistance is greatly appreciated

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

are they asking about the interest rate or just interest? because interest would just be 24*25-500 =600-500 =100

OpenStudy (anonymous):

Mathway teaches you step by step its not cheating

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

mathway isn't gonna help me for casas exam no calculators are allowed just raw math

OpenStudy (anonymous):

basically you are paying 4 dollars extra per payment